# Using iterated local alignment to aggregate trajectory data into a traffic flow map

### v2

@@ 5 Conclusion
We have introduced novel analysis algorithms to compute a flow map from empirical trajectories. Our starting point is to focus on aligning segments of the map matched routes rather than the complete routes. We define a spatial relation to set up local reference road segments, which allows us to align other nearby road segments to this local reference segment. This local alignment is the key innovation to computing a minimal flow map that is aligned to the underlying road network. We presented solid evidence for the high level of spatial resolution, accuracy and coverage for our proposed minimal flow map. Since it accurately shows the traffic flow on all road segments at all scales, it provides increased added value in comparison to the empirical trajectories, to the low resolution desire lines map, and to existing high resolution flow map methodologies.

### v5

@@ 4 Conclusion
We have introduced an iterative workflow which takes noisy GNSS vehicle trajectories as input and gives an accurate traffic flow map as output. Local alignment of flow segments is the key innovation. To carry out this local alignment, we defined a spatial relation to set up local reference flow segments, which then allowed us to align other nearby flow segments to this local reference segment. We presented solid evidence for the high level of spatial resolution and accuracy for our proposed locally aligned flow map for both synthetic and empirical trajectories. We demonstrated that they effectively reduce the noise of the input trajectories, and accurately aggregate the traffic flows on all road segments, and so they are a valuable addition to the toolkit for multi-scale traffic flow analysis.